Operations Report – November 15th 2018

Crew 198 Operations Report 15-Nov-2018

SOL: 5

Name of person filing report: Dana Levin

Non-nominal systems: Report filing, Rover Demos, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator update; failed overnight again, unable to obtain stable spin rates longer than 40 minutes. Todays attempt will be to start the generator at sunset regardless of SOC

– Report filing repeating from last night for broadcasting: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: functioning nominally

– Curiosity: functioning normally

– Opportunity: functioning normally

– Engineering air lock light and power sockets are still inoperable.

– Demos: rover remains inoperable.

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 11

Solar— SOC % (Before generator is run at night): 90%

Diesel Reading – 55%

Propane Reading – not taken

Ethanol Free Gasoline – 10 gallons.

Water (auxillary tank) – 80 gallons

Water (static tank) – 270 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 60

Water in GreenHab – tank at 40%

Water (loft) – Static to Loft Pump used – yes

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: no

Sojourner rover used: no

Hours: 2 cumulative

Beginning charge: 100%

Ending charge: 30

Currently charging: yes

Spirit rover used: no

Hours: 1

Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used: yes

Hours: 5

Beginning charge: 100

Ending charge: 21

Currently charging: yes

Curiosity rover used: yes

Hours: 5

Beginning charge: 100

Ending charge: 26

Currently charging: yes

Notes on rovers: (See above)

Demos: batteries refilled, charged at 100% but rover remains inoperable

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? Yes to obtain water to refill static tank

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional

Summary of suits and radios: using our own

Summary of Hab operations: nominal despite facility difficulties

Summary of GreenHab operations: none

Summary of ScienceDome operations: none

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

Operations Report Nov 14th

Crew 198 Operations Report 14-Nov-2018

SOL: 4

Name of person filing report: Dana Levin

Non-nominal systems: Report filing, Rover Demos, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator update; functioning nominally

– Report filing repeating from last night for broadcasting: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: functioning nominally

– Curiosity: battery died on EVA along Galileo road, these batteries appeared low on water as well. This rover will need to be retrieved

– Engineering air lock light and power sockets are still inoperable.

– Demos: battery refilled with water, appears to hold charge However rover remains inoperable.

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 11

Solar— SOC % (Before generator is run at night): 77%

Diesel Reading – 55%

Propane Reading – not taken

Ethanol Free Gasoline – 12 gallons.

Water (auxillary tank) – 60 gallons

Water (static tank) – 310 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 30

Water in GreenHab – tank at 75%, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 85

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 0

Currently charging: yes

Notes on rovers: (See above)

Demos: batteries refilled, charged at 100% but rover remains inoperable

Curiosity: abandoned on Galileo road with a dead battery

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? Yes to retrieve EVAers on corner of Cow dung and Gallileo road

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional

Summary of suits and radios: using our own

Summary of Hab operations: nominal despite facility difficulties

Summary of GreenHab operations: none

Summary of ScienceDome operations: none

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

Operations Report – November 13th

Crew 198 Operations Report 13-Nov-2018

SOL: 3

Name of person filing report: Dana Levin

Non-nominal systems: Report filing, Rover Demos, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator update; after resetting the inverters to enable it to communicate load requirements to the generator generator seems to be responding to load needs and functioned without difficulty last night. It was started at 83% today and appears to be functioning well as of this report

– Report filing repeating from last night for broadcasting: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: Shannon had used Sojourner two days ago but the battery drained to the point where she was unable to safely return to the habitat. Rover was left about 4 miles north of the Hab beside brae highway.

o Our schedule was rearranged to enable 2 of our instructors to leave the hab to recover the rover in the Hab car while our students were undergoing lectures by a third instructor. This was completed without incident and luckily our pre-planned schedule was sufficient to absorb this unexpected non crew/non sim related contingency.

– Engineering air lock light and power sockets are still inoperable.

– Demos Rover remains with battery drained, water has been replaced and rover is hooked up to charge cable, will evaluate in morning

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 12

Solar— SOC % (Before generator is run at night): 83% (see above)

Diesel Reading – 60%

Propane Reading – not taken

Ethanol Free Gasoline – 13 gallons.

Water (auxillary tank) – 64 gallons

Water (static tank) – 350 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 30

Water in GreenHab – tank at 90%, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge: 0%

Ending charge:

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 85

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 81

Currently charging: yes

Notes on rovers: (See above)

Demos: batteries refilled, attempting recharging overnight

Sojourner recovered and appears functional

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? No

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional
Summary of suits and radios: using our own
Summary of Hab operations: nominal despite facility difficulties Summary of GreenHab operations: none
Summary of ScienceDome operations: none
Summary of RAMM operations: none
Summary of any observatory issues: none identified as we are not using observatory
Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

operations report – november 12th

Crew 198 Operations Report 12-Nov-2018

SOL: 2

Name of person filing report: Dana Levin

Non-nominal systems: Generator, Report filing, Rover Sojourner, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator: Last night the generator had an underspeed cutout failure. While our crew has not been trained on generator mantinance and repair we were able to obtain assistance from our local benevolent Martian overlord Shannon.

– Event log;

o Around 1930 generator started with SOC at 46%

o 2000 generator cutout, temperature reading at 212F reporting a generator underspeed alert, generator restarted

o 2100 generator failed again, restarted Shannon alerted

o 2140 generator failed again and inverters cutout killing power to the entire MDRS complex. We drained coolant from the overflow tank including removing some congealed substance off the top of the tank, generator left off overnight

o At 0645 SOC was 26% solar began charging batteries

o 0800 generator started after evaluation by DG at 1800 RPM, continued operation for 1 hour and was shut off

o TODAY: 1720 SOC at 84% generator started as sunset. 1745 generator cutoff again.

– Current theory; mechanical obstruction to generator prevents full RPM which means current ops normal habitat power load exceeds generation capacity and generator stalls out

– Report filing: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: Shannon had used Sojourner two days ago but the battery drained to the point where she was unable to safely return to the habitat. Rover was left about 4 miles north of the Hab beside brae highway.

o Our schedule was rearranged to enable 2 of our instructors to leave the hab to recover the rover in the Hab car while our students were undergoing lectures by a third instructor. This was completed without incident and luckily our pre-planned schedule was sufficient to absorb this unexpected non crew/non sim related contingency.

– Engineering air lock light and power sockets are still inoperable.

– Demos Rover remains with battery drained and empty of water, rover is inoperable.

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 2

Solar— SOC % (Before generator is run at night): 84% (see above)

Diesel Reading – 60%

Propane Reading – not taken

Ethanol Free Gasoline – 14 gallons.

Water (auxillary tank) – 60 gallons

Water (static tank) – 380 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 20

Water in GreenHab – full tank, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: recovered and currently charging

Hours: 1 hour

Beginning charge: 31%

Ending charge: 40% (towed)

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 89

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 81

Currently charging: yes

Notes on rovers: (See above)

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? Yes, to retrieve sojourner from Brae highway and to pump water into greenhab

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional

Summary of suits and radios: using our own

Summary of Hab operations: nominal despite facility difficulties

Summary of GreenHab operations: none

Summary of ScienceDome operations: none

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

Operations Report – November 11th 2018

Crew 198 Operations Report 11/10/2018

SOL: 1

Name of person filing report: Dana Levin

Non-nominal systems: Engineering Lock light and power plugs near back section of habitat

Notes on non-nominal systems: they’ve been non functional since our first mission 5 years ago at this point it may be more accurate to refer to the non functional state as “nominal” I will let you know if they suddenly become “off-nominal” and turn on.

Generator (hours run): 7 hours (midnight to 0700)

Solar— SOC 65% (Before generator is run at night)

Diesel – 60%

Propane – psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 15 gallons.

Water (trailer) – 400 gallons.

Water (static) – 60 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter:

ATV’s Used: (Honda, 350.1, 350.2)

Oil Added? No

ATV Fuel Used: 0.5 Gals each

# Hours the ATVs were Used today: .75h

Notes on ATVs: HOnda ATV will not shift into reverse

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: disabled

Hours: 0.75

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Spirit rover used: yes

Hours: 0.75

Beginning charge: 100

Ending charge: 89

Currently charging: yes

Opportunity rover used: yes

Hours: 0.75

Beginning charge: 100

Ending charge: 88

Currently charging: yes

Curiosity rover used: yes

HabCar used and why, where? Yes to pick up water from hollow mountain (2 runs)

General notes and comments:

Summary of internet: functional

Summary of suits and radios: suits functional, radios not used/ 2MARS crews use our own comms

Summary of Hab operations: today was orientation and medical scenarios

Summary of GreenHab operations: maintenance

Summary of ScienceDome operations: maintenance and engineering

Summary of RAMM operations: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: none at present.

Operations report – October 25th

Crew 197 Operations Report 25Oct18
SOL: 4
Name of person filing report: Jim Ehrhart
Non-nominal systems: solar/batteries
Notes on non-nominal systems: batteries apparently on site, to be installed Friday
Generator (hours run): 24hrs. The battery to start the generator is sitting on the ground. This can lead to depletion is not charging from the generator. Will put on wood blocks tomorrow.
Solar— SOC % (Before generator is run at night): N/A
Diesel: 65-70%, hard to read precisely from ground due to parallax Propane – psi: 45
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons: Not checked, ATV not in use
Water (trailer) – gallons: N/A trailer not in use
Water (static) – gallons: estimated 75-80% full, no precise gauge available Trailer to Static Pump used – no or yes: No
Water (loft) – Static to Loft Pump used – yes or no: Yes
Water Meter: morning fill, from 25 to 50, meter 138088.98; evening fill from 25 to 50, meter 138160.28
Toilet tank emptied: no or yes: No
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Oil Added? No
ATV Fuel Used: Gals No
# Hours the ATVs were Used today: 0
Notes on ATVs: N/A
Deimos rover used: No, out of service
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unk
Beginning charge: Unk
Ending charge: Unk
Currently charging: Unk
Spirit rover used: Yes
Hours: 53.9 hrs
Beginning charge: 100
Ending charge: 60%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 37.0 hrs
Beginning charge: 100
Ending charge: 41%
Currently charging: No
Opportunity experienced non-nominal behavior during EVA 3. Rover was carrying larger team members, Robert (5’7”) and James (6’9”) and their knees apparently hit the gear buttons randomly that resulted in the rover being in High gear for portions of the trip and rapid battery consumption. Approximately 100m from the Hab the rover was at 25% and the “motor overheat” lamp came on and the rover stopped. The crew pushed it back to its proper parking spot and used a spare extension cord from the GreenHab to plug it in.
The Director subsequently unplugged it so that only one rover (Curiosity) was plugged in to the existing cord. (Spirit is plugged in at the GreenHab). At the time of the Engineering EVA (5:30pm) Curiosity had reached 41% (prior to being unplugged) and the “motor overheat” button no longer came on. No test drive was performed. Curiosity rover used: Yes
Hours: 52.3 hrs
Beginning charge: 100
Ending charge: 86%
Currently charging: Yes
HabCar used and why, where? No
General notes and comments:
Summary of internet: Internet connectivity and bandwidth fine Summary of suits and radios: All suits performed nominally. One charger missing and one wonky – attempted to create two new chargers from existing spare parts with suit engineers, but couldn’t produce working unit from spares. Need to send to NorCal chapter for replacement. Suits experienced wear and tear on EVAs, Suit 9 got a significant scratch on helmet
All radios reviewed and settings standardized. 13 units in good shape and fully charged. 4 units not in good condition and set aside as reserve.
Summary of Hab operations: All systems nominal.
Summary of GreenHab operations: Nominal, smoke detector batteries replaced and tested. Plants sprouted.
Summary of ScienceDome operations: N/A, not in use due to battery issue Summary of RAMM operations: Nominal, very disorderly, needs cleaning and organization
Summary of health and safety issues: Health and Safety Officer perform checks – see report.
Questions, concerns and requests to Mission Support:
– Battery meter acquired as donation from Shannon Norrrel. Will put in RAMM for use in checking status of battery spares.

Operations Report Oct 24th

  • Crew 197 Operations Report 24Oct18
    SOL: 3
    Name of person filing report: Jim Ehrhart

Non-nominal systems: solar/batteries
Notes on non-nominal systems: batteries apparently on site, to be
installed Friday
Generator (hours run): 23.5, generator on almost full-time, down 5 min
at approx. 5:30pm to add coolant
Solar— SOC % (Before generator is run at night): N/A
Diesel: 75%
Propane – psi: 45
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons: Not
checked, ATV not in use

Water (trailer) – gallons: N/A trailer not in use
Water (static) – gallons: estimated 90% full
Trailer to Static Pump used – no or yes: No
Water (loft) – Static to Loft Pump used – yes or no: Yes
Water Meter: morning fill, from 10 to 50, meter 138088.98; evening
fill from 10 to 50, meter 138112.92
Toilet tank emptied: no or yes: Yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Oil Added? No
ATV Fuel Used: Gals No
# Hours the ATVs were Used today: 0
Notes on ATVs: N/A

Deimos rover used: No, out of service
Hours: 134
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unk
Beginning charge: Unk
Ending charge: Unk
Currently charging: Unk
Spirit rover used: Yes
Hours: 42.3
Beginning charge: 100
Ending charge: 99
Currently charging: Yes
Opportunity rover used: Yes
Hours: 36.1 hrs
Beginning charge: 100
Ending charge: 100
Currently charging: No
Curiosity rover used: Yes
Hours: 55.3 hrs
Beginning charge: 100
Ending charge: 97
Currently charging: Yes

HabCar used and why, where? No

General notes and comments:

Summary of internet: Internet connectivity and bandwidth fine
Summary of suits and radios: All suits performed nominally. One
charger missing and one wonky – will create two new chargers from
existing spare parts tomorrow with suit engineers. Suits experienced
wear and tear on EVAs, some quilting torn in rover ingress/egress, two
scratches on one helmet
All radios reviewed and settings standardized. 13 units in good shape
and fully charged. 4 units not in good condition and set aside as
reserve.
Summary of Hab operations: All systems nominal. Fan/heater air filter
replaced. Loft water filter replaced.
Summary of GreenHab operations: Nominal, smoke detector batteries
replaced and tested
Summary of ScienceDome operations: N/A, not in use due to battery issue
Summary of RAMM operations: Nominal, very disorderly, needs cleaning
and organization
Summary of health and safety issues: Health and Safety Officer perform
checks – see report.
Questions, concerns and requests to Mission Support:
– Battery meter needed to check random spares found in RAMM.
– Suggest light in loft

Operations Report October 22nd

Operations Report 22OCT2018

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Batteries for power system

Notes on non-nominal systems: Batteries arrive tomorrow and will be installed on Friday!

Generator (hours run): 24

Solar— SOC n/a %

Diesel – 80 %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (axillary tank) – 0 gallons.

Water (static) – 500 gallons

GreenHab water tank: 200 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda used to move Deimos into charging position, the 350 that was not starting needed fuel, because I switched to the reserve tank and it worked fine. This is interesting, because it really did not look empty, it looked about 1/3 fuel. All ATVs, if used, should have a full tank when leaving MDRS. The three 350’s and the 300 have been taken out of sim service. The Honda remains, but only to be used if a rover breaks down and needs tows back to MDRS. Crews will not be trained on the use of ATVs, but one experienced person or chosen person per crew will be trained on how to tow.

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs:

Deimos rover used: Yes, towed into position in the charging bay next to Hab. I will work on it there,

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used: yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: no.

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: no

HabCar used and why, where? Was used in the am to haul water, then driven to Bicknell for the following repairs: New windshield, new taillight, oil change and safety inspection. We will pick it up at the end of the week and get new tabs for it.

Crew Transport Vehicle: Yes, to drop the HabCar off and to get fence posts for the tunnel

General notes and comments: Work on the tunnel was started today but not finished due to (you guessed it!) rain. Most of the fenceposts were reset and four of the tarps are in place.

Summary of internet: All three accounts nominal. HAL is
operational. More information will be coming about this system.

Summary of suits and radios: Suits have arrived and are being charged for training in the morning.

Summary of Hab operations: All materials have been removed from in front of the Hab except the rocks, which we are referring to as “the in situ resource material”. I will move it over time.

There is a leak in the drain pipe from the summer work done on the plumbing. It is where the old science sink used to drain. For now, we are using the upstairs sink normally, and catching the water in a bucket using the contractor can fix it.

We will be turning on the hot water in the morning and conducting a propane test in the morning.

Summary of GreenHab operations: We began using the big blue water tank in the GreenHab. It is nice to be able to water from inside the building and will provide some temperature buffering capacity.

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of health and safety issues: nothing to report.

Notes, questions, concerns and requests to Mission Support: nothing at this time

Operations Report – June 4th

SOL: 0

Name of person filing report: Shannon Rupert

Non-nominal systems: Do you believe in magic? Because at this very moment, we have no non-nominal systems! (Knock wood)

Notes on non-nominal systems: Happy dance!

Generator (hours run): Turned off at 8 am, will turn back on at 8 pm

Solar—100 % SOC

Diesel – 50 %

Propane – 58 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –Not checked

Water (trailer) –450 gallons.

Water (static) – 450 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: yes, and flushed twice by filling and emptying.

ATV’s Used: Yes

Oil Added? no

ATV Fuel Used: n/a

# Hours the ATVs were Used today: 1

Notes on ATVs: Mostly moving around, some training. Need new bolts for new battery for 300.

Deimos rover used: Yes

Hours:134.0

Beginning charge: 100

Ending charge:100

Currently charging: no

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 42.3

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 31.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes (NOTE: Driver’s seat is bent, should be replayed by one from Phobos)

Curiosity rover used: yes

Hours: 39.0

Beginning charge: 97

Ending charge: still being used

Currently charging: No

HabCar used and why, where? Yes, to get water and move things from MDRS to town, also trash

General notes and comments: The Korean Project began last evening as the first of the team arrived. We did some generally cleaning. This morning the transformation began. We have 40 plus people on site. The Hab was given some snazzy upgrades in terms of lower deck tables and upper deck kitchen. An airstream was brought in to serve as their production center and is parked on the opposite side of the GreenHab and tunnel. Four motorhomes, a bathroom trailer, water truck and large generator are now parked in my yard. Tomorrow they are putting in a misting shelter. I am looking forward to this as I can’t use my AC due to their power needs.

Cameras have been installed everywhere. Portable toilets have all been removed from campus and URC sites.

Summary of internet: Nothing to report

Summary of suits and radios: The crew installing small cameras on the new faceplates and understand that if there is any damage they will be charged $50 per faceplate.

Summary of Hab operations: The AC (swamp cooler) has been installed on the upper deck and it was working wonderfully until the crew emptied the water tank. I got it up and running again but not sure if the water is working correctly. I will check later.

Summary of GreenHab operations: 50 plus new plants growing food will be installed in the GreenHab this evening for use by the crew for food. Plants were grown off station for this project,.

Summary of ScienceDome operations: Lab fridge is clean, both fridge and freezer were relocated to the lab table and off the floor. New lamp left by the WMS crews was installed and looks great.

Summary of RAM operations: The tool box on the lower deck was relocated to its new home in the RAM. We have cleaned the RAM for the filming but it is not organized properly.

Summary of health and safety issues: Because so much trash was generated by the URC, we burned it and now have a bit of a mess to clean up. Broken and burned glass and metal needs to be removed and taken to town.

The road is so sand pitted now that we had many people stuck today.

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – May 31st

SOL: 31May2018

Name of person filing report: Shannon Rupert

Non-nominal systems: Only me! Only me!

Notes on non-nominal systems: This morning I took on the power system again. I changed two settings on the AC controls to what I believed meant the solar took precedence over the generator for delivery of power to the generator. I had been working on this system without any improvement for so long that I didn’t give it another thought except that I would have to keep running the generator as the SOC fell during the day. But it never fell! We are back on solar power. I’m not declaring complete victory yet, but I am hopeful. In addition, the contractor will be here on Saturday to go over the system with me.

Generator (hours run): Turned off at 8 am, will turn back on at 8 pm

Solar—100 % SOC

Diesel – not checked %

Propane – not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –Not checked

Water (trailer) – 0 gallons.

Water (static) – not checked gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: yes. I intended to flush it again and have no one use it until the tv crew arrives. But people used it anyway. So I have to do it again. Sigh.

ATV’s Used: Yes

Oil Added? Yes

ATV Fuel Used: n/a

# Hours the ATVs were Used today: 1

Notes on ATVs: We just used them to block the driveway today so people did not drive up to the Hab.

Deimos rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? No

General notes and comments: AC to GreenHab completed. AC in Hab not yet installed.

Today was the first full day of the URC and my judging duties kept me from working on the things I had hoped to do. They will have to wait, too tired tonight to do anything more.

Summary of internet: Nothing to report

Summary of suits and radios: Suits in RAM. Six remaining new radios put into service.

Summary of Hab operations: Chaos reigns.

Summary of GreenHab operations: Water was hooked up to the swamp cooler so we now have AC in the greenhouse. URC operations on lower deck.

Summary of ScienceDome operations: URC operations happened today. We discovered that the fridge in the lab is filled with moldy petri dishes. I will clean it out in the morning before we start the URC events,

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time